3.40*10-11 = ( 3.00*10-3) X2

1
Expert's answer
2021-05-25T07:50:41-0400

3.40 × 10-11 = ( 3.00 × 10-3) X2

X2 = 3.40 × 10-11 / 3.00 × 10-3

X2 = 1.13 × 10-8

X = (1.13 × 10-8)1/2 = 1.06 × 10-4


Answer: 1.06 × 10-4